# On automatic boundedness of operators between ordered and topological vector spaces

### 11 Mar 2026 v2

@@ 2. $b)$
    $(\sigma$-))order continuous if $x_{n}\operatorname{\xrightarrow[]{\tau}}0$ (resp., $\|x_{n}\|\to 0$) for every sequence $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$.
@@
Let $X$ be an OBS with a closed normal cone.
@@ 1. $i)$
    If the norm in $X$ is order continuous then $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$ in $X$.
@@
    If the norm in $X$ is $\sigma$-order continuous then $x_{n}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{n}\operatorname{\xrightarrow[]{(ru)}}0$ in $X$.
@@
First, we show that if $X_{+}$ is normal then $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0\Longrightarrow x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Suppose $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$. Then, for some $u\in X_{+}$ there exists an increasing sequence $(\alpha_{n})$, $\frac{1}{n}u\leq x_{\alpha}\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$. Let $0\leq y_{0}\leq\frac{1}{n}u$ for all $n\in\mathbb{N}$. Since $\big\|\frac{1}{n}u\big\|\to 0$ then $y_{0}=0$ by [4, Theorem 2.23]. It follows $\frac{1}{n}u\downarrow 0$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$.
$i)$ Let $X$ has order continuous norm and $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Take a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is $\alpha_{\beta}$ satisfying $\pm x_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$. Find an increasing sequence $(\beta_{k})$ such that $\sup\limits_{z\in[-g_{\beta_{k}},g_{\beta_{k}}]}\|z\|\leq k^{-3}$ for all $k\in\mathbb{N}$, and set $u=\|\cdot\|$-$\sum_{k=1}^{\infty}kg_{\beta_{k}}$. Then $\pm x_{\alpha}\leq g_{\beta_{k}}\leq k^{-1}u$ for $\alpha\geq\alpha_{\beta_{k}}$. Choose any increasing sequence $(\alpha_{k})$ with $\alpha_{k}\geq\alpha_{\beta_{1}},\alpha_{\beta_{2}},\ldots,\alpha_{\beta_{k}}$. Then $\pm x_{\alpha}\leq k^{-1}u$ for $\alpha\geq\alpha_{k}$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{(ru)}}0$.
@@
Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS.
@@
We prove only $i)$ as the proof of ii)ii) is similar. In view of Lemma 3.8, we need to prove ${\cal L}^{\sigma}_{Leb}(X,Y)\subseteq{\cal L}^{\sigma}_{onc}(X,Y)$ and ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y)$. Let $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$. Then $x_{n}\operatorname{\xrightarrow[]{(ru)}}0$, by Lemma 3.15, and hence $\|x_{n}\|\to 0$ due to normality of $X_{+}$.
If $T\in{\cal L}^{\sigma}_{Leb}(X,Y)$ then $T\in{\rm L}(X,Y)$ by Theorem 3.11, and hence $\|Tx_{n}\|\to 0$. If $T\in{\cal L}^{\sigma}_{wLeb}(X,Y)$. Then again $T\in{\rm L}(X,Y)$ by Theorem 3.11. So $\|Tx_{n}\|\to 0$ and $Tx_{n}\operatorname{\xrightarrow[]{w}}0$. ∎
The next corollary is a partial extension of results of [22, Lemma 1], [5, Lemma 2.1], [25, Theorem 3], and [23, Proposition 1.5] to the OVS setting.
@@
Let $E$ be a Banach lattice with order continuous norm and $Y$ be an NS. Then ${\cal L}^{\sigma}_{Leb}(E,Y)={\cal L}^{\sigma}_{onc}(E,Y)$, ${\cal L}_{Leb}(E,Y)={\cal L}_{onc}(E,Y)$, ${\cal L}^{\sigma}_{wLeb}(E,Y)={\cal L}^{\sigma}_{owc}(E,Y)$, and
${\cal L}_{wLeb}(E,Y)={\cal L}_{owc}(E,Y)$.
@@
Let $X$ be an OBS with a closed normal cone and $Y$ be an ONS with a generating normal cone.
@@
We prove only $i)$ as the proof of ii)ii) is similar. Let $T\in{\cal L}_{ob}(X,Y)$ and $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$. By Lemma 3.15, $x_{n}\operatorname{\xrightarrow[]{(ru)}}0$. It follows from [16, Corollary 2.5] that $Tx_{n}\operatorname{\xrightarrow[]{(ru)}}0$. The normality of the cone $Y_{+}$ provides $\|Tx_{n}\|\to 0$. ∎

### 16 Apr 2026 v3

@@ 2. $b)$
    $\sigma$-order continuous if $x_{n}\operatorname{\xrightarrow[]{\tau}}0$ (resp., $\|x_{n}\|\to 0$) for every sequence $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$.
@@
Let $X$ be an OBS with a closed normal cone. The following assertions hold.
@@ 1. $i)$
    If the norm in $X$ is order continuous then $x_{\alpha}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$ in $X$.
@@
    If the norm in $X$ is $\sigma$-order continuous then $x_{n}\operatorname{\xrightarrow[]{o}}0\Longleftrightarrow x_{n}\operatorname{\xrightarrow[]{ru}}0$ in $X$.
@@
First, we show that if $X_{+}$ is normal then $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0\Longrightarrow x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Suppose $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$. Then, for some $u\in X_{+}$ there exists an increasing sequence $(\alpha_{n})$ with $\frac{1}{n}u\leq x_{\alpha}\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$. Let $0\leq y_{0}\leq\frac{1}{n}u$ for all $n\in\mathbb{N}$. Since $\big\|\frac{1}{n}u\big\|\to 0$ then $y_{0}=0$ by [4, Theorem 2.23]. It follows $\frac{1}{n}u\downarrow 0$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$.
$i)$ Let $X$ has order continuous norm and $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. Take a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is $\alpha_{\beta}$ satisfying $\pm x_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$. Find an increasing sequence $(\beta_{k})$ such that $\sup\limits_{z\in[-g_{\beta_{k}},g_{\beta_{k}}]}\|z\|\leq k^{-3}$ for all $k\in\mathbb{N}$, and set $u=\|\cdot\|$-$\sum_{k=1}^{\infty}kg_{\beta_{k}}$. Then $\pm x_{\alpha}\leq g_{\beta_{k}}\leq k^{-1}u$ for $\alpha\geq\alpha_{\beta_{k}}$. Choose any increasing sequence $(\alpha_{k})$ with $\alpha_{k}\geq\alpha_{\beta_{1}},\alpha_{\beta_{2}},\ldots,\alpha_{\beta_{k}}$. Then $\pm x_{\alpha}\leq k^{-1}u$ for $\alpha\geq\alpha_{k}$, and hence $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$.
@@
Let $X$ be an ordered Banach space with a closed generating normal cone and $Y$ a normed space. The following assertions hold.
@@
We prove only $i)$ as the proof of ii)ii) is similar. In view of Lemma 3.9, we need to prove ${\cal L}^{\sigma}_{Leb}(X,Y)\subseteq{\cal L}^{\sigma}_{onc}(X,Y)$ and ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y)$. Let $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$. Then $x_{n}\operatorname{\xrightarrow[]{ru}}0$, by Lemma 3.16, and hence $\|x_{n}\|\to 0$ due to normality of $X_{+}$.
If $T\in{\cal L}^{\sigma}_{Leb}(X,Y)$ then $T\in{\cal L}_{b}(X,Y)$ by Theorem 3.12, and hence $\|Tx_{n}\|\to 0$. If $T\in{\cal L}^{\sigma}_{wLeb}(X,Y)$. Then again $T\in{\cal L}_{b}(X,Y)$ by Theorem 3.12. So $\|Tx_{n}\|\to 0$ and $Tx_{n}\operatorname{\xrightarrow[]{w}}0$. ∎
The next corollary is a partial extension of results of [21, Lemma 1], [5, Lemma 2.1], [24, Theorem 3], and [22, Proposition 1.5] to the OVS setting.
@@
Let $E$ be a Banach lattice with order continuous norm and $Y$ a normed space. Then ${\cal L}^{\sigma}_{Leb}(E,Y)={\cal L}^{\sigma}_{onc}(E,Y)$, ${\cal L}_{Leb}(E,Y)={\cal L}_{onc}(E,Y)$, ${\cal L}^{\sigma}_{wLeb}(E,Y)={\cal L}^{\sigma}_{owc}(E,Y)$, and ${\cal L}_{wLeb}(E,Y)={\cal L}_{owc}(E,Y)$.
@@
Let $X$ be an ordered Banach space with a closed normal cone and $Y$ an ordered normed space with a generating normal cone. The following assertions hold.
@@
We prove only $i)$ as the proof of ii)ii) is similar. Let $T\in{\cal L}_{ob}(X,Y)$ and $x_{n}\operatorname{\xrightarrow[]{o}}0$ in $X$. By Lemma 3.16, $x_{n}\operatorname{\xrightarrow[]{ru}}0$. It follows from [15, Corollary 2.5] that $Tx_{n}\operatorname{\xrightarrow[]{ru}}0$. The normality of the cone $Y_{+}$ ensures $\|Tx_{n}\|\to 0$. ∎